OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x - 2y = 8.

OpenStudy (anonymous):

x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (-8, 0); y intercept = (0, -4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (anonymous):

-2x=-x+8

OpenStudy (alexwee123):

sub in 0 for y then solve for x to find x intercept x-2(0)=8 x=8 then do the opposite to find the y int 0-2y=8 y=-4 so you intercepts are (8,0) and (0,-4)

OpenStudy (anonymous):

y=1/2x-4 is slope intercept form

OpenStudy (alexwee123):

so it's the second option